# Long-Context Autoregressive Video Modeling with Next-Frame Prediction

### v1

@@ 5.4 Ablation Study
Stochastic Clean Context. We have visualized the effectiveness of stochastic clean context in Fig. 4. Based on the quantitative evaluation of video prediction in Tab. 5, FAR with stochastic clean context achieves significantly improved performance.
Table 5: Ablation Study of Stochastic Clean Context on UCF-101. Stochastic clean context mitigates the training-inference discrepancy in observed context, leading to improved performance.
@@
Table 6: Ablation Study on the Resolution of Long-Term Context. The speed is averaged over the generation of 300 frames.
@@ 6 Conclusion
In this paper, we introduce FAR for autoregressive video modeling. FAR enables learning causal dependency of continuous frames and demonstrates better convergence than video diffusion transformers. Building upon FAR, we explore its application in long-context video modeling. We identify visual redundancy as a key challenge, leading to poor temporal extrapolation performance at test time and inefficient long-video training. To address these issues, we propose balancing locality and long-range dependency. Specifically, we introduce FlexRoPE to enable 16$\times$ temporal extrapolation at test time and incorporate long short-term context modeling for efficient long-video training. FAR achieves state-of-the-art performance on both short- and long-video modeling, highlighting its potential as a new foundation model for video generation.

### v2

@@ 5.4 Ablation Study
Stochastic Clean Context. We have visualized the effectiveness of stochastic clean context in Fig. 4. Based on the quantitative evaluation of video prediction in Table. VI, FAR with stochastic clean context achieves significantly improved performance.
Multi-Level KV Cache. We evaluate the inference speed of FAR in Fig. 9. FAR without the KV cache incurs the highest inference cost, while the KV cache significantly reduces it. For long videos, FAR-Long with proposed multi-level KV cache achieves the lowest inference cost.
TABLE VI: Ablation Study of Stochastic Clean Context on UCF-101. Stochastic clean context mitigates the training-inference discrepancy in observed context, leading to improved performance.
@@
TABLE VII: Ablation Study on the Resolution of Long-Term Context. The speed is averaged over the generation of 300 frames.
@@ 6 Conclusion
In this paper, we introduce FAR for long-context video modeling. FAR effectively learns causal dependencies across continuous frames and demonstrates faster convergence compared to video diffusion transformers. To address visual redundancy in long-video training, we further propose a long short-term context modeling strategy, which encodes both remote and nearby visual contexts using asymmetric patchification. This approach enables a manageable token length as context frames increase, reducing both time and memory costs during long-video training. FAR achieves state-of-the-art performance on both short- and long-video modeling tasks, underscoring its potential as a new foundation model for video generation.
Limitations. The primary limitation lies in the lack of scaled-up experiments. Although FAR demonstrates great potential, we still lack large-scale training on text-to-video generation datasets. Additionally, restricted by the available datasets, we only experiment with FAR on up to 300 frames (about 20 seconds), not fully investigating its ability on minute-level videos.
Future Work. One future direction is to scale up FAR and benchmark it against video diffusion transformers on large-scale text-to-video generation tasks. Additionally, we plan to simulate a longer video dataset (on the minute level) to better evaluate the model’s long-context capabilities. Finally, it would be interesting to explore whether FAR’s long-context modeling can enable video-level in-context learning.